OpenStudy (anonymous):

Can you help me with writing each ratio as a fraction in simplest form??? I'm sorry i won't reply to your. Comments i don't know how to send my comments! 2. 12 to 8 4. 6:50 6. 3:8 8. 8 out of 11

OpenStudy (anonymous):

2. \[\frac{12}{8} = \frac{3}{2}\] 4. \[\frac{6}{50} = \frac{3}{25}\] 6. \[\frac{3}{8}\] 8. \[\frac{8}{11}\]

OpenStudy (mathmale):

Pinky: At the bottom of each screen like this one is a box in which you see "Type your answer here" or something like that. Do that, and then click on the Post button (below, right). In 12/8, linh correctly recognized that 4 is common to both 12 and 8, so he factored out the 4 and was left with \[\frac{ 12 }{ 8 }=\frac{ 4(3) }{ 4(2)}.\] Then he cancelled out the 4, which again was correct.

OpenStudy (mathmale):

Pinky, please try that yourself. How would you explain the solution of the next problem on your list?